python Hi guys, I really need help with that please Can I get help with this...

python

Hi guys, I really need help with that please

Can I get help with this python program?

Please write clear

Show all the steps please, the outcome and follow the directions.

Volume
# Purpose: This program computes the volume (in liters) of a six-pack of soda cans and the total volume of a six-pack and a two-liter bottle.

Complete the following:

# Liters in a 12-ounce can and a two-liter bottle.


# Number of cans per pack.


# Calculate total volume in the cans and print the statement.


# Calculate total volume in the cans and a 2-liter bottle; and print the statement.
Answer #1

Explanation :

It is given six-pack soda cans so we know that total number of cans is 6 and it is also given that the volume of one can is 12 ounce so we convert 12 ounce to liter and as 1 ounce=0.0295735 liter so as 12 ounce.

and then add 2 to get the net volume.

CODE:

# first of all convert ounces to liters and compute the volume of one can into liters because its given that the volume of 1 can is 12 ounces
x=12
# volume of a can two lter bottle
vol_can_lit=x*0.0295735
# volume of a can and two lter bottle
vol_can2_lit=x*0.0295735+2
# number of cans in a pack
y=6
#total volume in cans
net_vol_cans=12*vol_can_lit
print("Total volume in cans:",net_vol_cans)
print("Total volume in cans and a 2-liter bottle",vol_can2_lit)